### v1

@@ 4.2 Arbitrary groups
We now allow the group to be arbitrary. The classic Bahadur & Savage (1956) result establishes the impossibility of testing null hypotheses that are too ‘dense’ within the space of all continuously distributed distributions. In the context of analyzing randomization tests, it turns out that a similar conclusion holds under a related local version of this denseness property. Let $\Omega_{\text{cts}}(\mathcal{B})$ be the set of distributions supported on $\mathcal{B}\subseteq\mathbb{R}$.
@@
A set $\Omega_{0}\subset\Omega_{\text{cts}}(\mathcal{B})$ is locally-dense if, roughly speaking, any density on a finite union of small intervals can be matched (up to scale) by some density in $\Omega_{0}$. Formally, it is locally-dense if there exists $L:\mathbb{N}\to\mathbb{R}_{+}$ such that for any intervals $I_{1},\dots,I_{m}$ with $|I_{i}|<L(m)$ and any density $p$ supported on these intervals, there exists $H\in\Omega_{0}$ with density $h$ and scaling $\alpha\in(0,1]$ satisfying $h(x)=\alpha p(x)$ on $\bigcup_{j=1}^{m}I_{j}$.
@@
The proof of Proposition 6 requires developing additional notation to address technical considerations arising from working with continuous distributions. The proof makes use of Proposition 1 and shows that for any function $f$ such that $\Omega_{0}\subseteq\Omega_{f}$, local denseness implies that $\Omega_{f}=\Omega$. Intuitively, the local denseness property ensures that the null hypothesis is locally rich enough to represent any density at finitely many locations.
@@
To provide some intuition for this result, consider the null hypothesis that $\operatorname{\mathbb{E}}_{P}[X_{i}]=\beta$. To show this null hypothesis is locally-dense, consider any set of $m$ intervals with width bounded by some $\delta>0$ (so that $L(m)=\delta$). For any density $p$ on these intervals with mean $\gamma_{p}$ and $\alpha\in(0,1]$, a distribution $H$ on the reals with density $\alpha p(x)$ on the intervals will have mean $\alpha\gamma_{p}+(1-\alpha)\operatorname{\mathbb{E}}_{H}[X_{i}|\text{not in intervals}]$. Since there always are two additional intervals that respectively lie above and below $\beta$ that do not intersect with the $m$ intervals, we can judiciously choose the weighting $\alpha$ and the mass on these two intervals so that the resulting $H$ has mean $\beta$. This shows that the null hypothesis is locally-dense. The following results all make use of similar constructions.
@@
To further illustrate the use of Proposition 6, we show in Supplementary Appendix B.3 that the statement in Proposition 7 holds if we replace $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and $\beta$ in the support. The same holds with Proposition 8 for any $q\in(b_{0},b_{1})$. In this appendix, we also show that the null hypothesis of the variance equaling some value is locally-dense and does not satisfy the randomization hypothesis.
@@ 5 Conclusion
Randomization tests yield exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. However, many randomization tests used in practice only provide such guarantees for nulls substantially stronger than the intended null. This paper developed a framework for characterizing which null hypotheses admit randomization tests with exact finite-sample Type 1 error control.
Applying our framework to one-sample problems revealed impossibilities and limited possibilities of uses of finite-sample randomization tests. Many standard null hypotheses, including the null of mean zero, cannot be tested using any randomization test while maintaining finite-sample validity. We further showed that linear group actions — the basis of all existing randomization tests — can only test subsets of symmetric or normal distributions in one-sample settings. These impossibility results confirm that practitioners using existing tests are not missing better alternatives. While our results establish that testing beyond symmetry or normality requires non-linear transformations, we also show that even such transformations cannot be used to test many common null hypotheses. The framework we developed not only delineates these boundaries but also provides constructive guidance for identifying when nulls admit randomization tests and explicitly building such tests when they exist. Overall, our findings validate the current practice of using standard randomization tests and further motivate focusing on the asymptotic validity of randomization tests.
@@ Appendix A Examples of randomization tests
    For $L>0$, let $\mathcal{A}\equiv\{-L,\dots,0,\dots,L\}$ and let $\Omega_{0}$ be the set of distributions that are symmetric around 0. Consider any $x,y$ with $x_{j}\neq 0$ for some $j$ and y=xy=-x. We then have that $c(x_{j};x)>c(x_{j};y)$ and that for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(-x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{\text{discrete}}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that send some values of $x\in\mathcal{A}^{n}$ to x-x and otherwise act as the identity function.
@@
    For any $\mathcal{A}$, let $\Omega_{0}$ be the set of distributions where $p(\alpha_{\ell})=p(\alpha_{\ell^{\prime}})$ for some $\alpha_{\ell}\neq\alpha_{\ell^{\prime}}$. Consider any $x,y$ with $x$ containing $\alpha_{\ell}$ but not $\alpha_{\ell^{\prime}}$ and $y_{j}=\alpha_{\ell^{\prime}}\mathbb{1}[x_{j}=\alpha_{\ell}]+\alpha_{\ell}\mathbb{1}[x_{j}=\alpha_{\ell^{\prime}}]+x_{j}\mathbb{1}[x_{j}\neq\alpha_{\ell},x_{j}\neq\alpha_{\ell^{\prime}}]$. We then have that $c(\alpha_{\ell};x)>c(\alpha_{\ell};y)$ and for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{\text{discrete}}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that acts as the identity except for swapping the values of $\alpha_{\ell}$ and $\alpha_{\ell^{\prime}}$.
@@
    Let $\mathbb{G}$ be the orthogonal group of degree $n$. By Maxwell’s Theorem, the set of invariances given $\mathbb{G}$ is $\Omega_{\mathbb{G}}=\{N(0,\sigma^{2}):\sigma\in\mathbb{R}\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian with mean zero.
@@
    In the below proof, we show that $\mathbb{G}$ is a group and that $\Omega_{\mathbb{G}}=\{N(\mu,\sigma^{2}):(\mu,\sigma)\in\mathbb{R}^{2}\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian.
@@
Let $X_{1},X_{2},X_{3}$ be i.i.d. draws from $N(\mu,\sigma^{2})$ for some $\mu,\sigma\in\operatorname{\mathbb{R}}$. Let $X=(X_{1},X_{2},X_{3})$ and $V=(\mu,\mu,\mu)$. Because
@@
Inductively, $A^{n}X\stackrel{{\scriptstyle d}}{{=}}X$ for all $n\in\mathbb{Z}$. Thus, $\{N(\mu,\sigma^{2}):\mu,\sigma\in\operatorname{\mathbb{R}}\}\subset\Omega_{\mathbb{G}}$. By proposition 5, $\{N(\mu,\sigma^{2}):\mu,\sigma\in\operatorname{\mathbb{R}}\}=\Omega_{\mathbb{G}}$. ∎
@@ B.1 All proofs excluding the proof of Proposition 6
$\Omega_{f}=\bigcap_{x\in\mathcal{A}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})\ \text{s.t.}\ y=f(x)\right\}.$ (10)
Then any $f$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$ must have some $x^{\star}\in\mathcal{A}$ and $y^{\star}=f(x^{\star})$ such that
@@
Let $x,y\in\mathcal{A}^{n}$ be arbitrary such that $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$ and let $\beta_{i}(x,y)=c(\alpha_{i};x)-c(\alpha_{i};y)$. We know that there exists some $1\leq i\leq K-1$ such that $\alpha_{i}\leq q<\alpha_{i+1}$. One of the follow two cases must occur:
@@
Since $\mathbb{G}$ is a group of linear maps, any element of the group is a bijective and thus invertible $n\times n$ matrix $A=\{a_{ij}\}_{ij}$. Suppose there exists $i,i^{\prime}$ and some $j^{\prime}$ such that $a_{ij^{\prime}},a_{i^{\prime}j^{\prime}}\neq 0$. If $\Omega_{A}$ is not empty, then
@@
where $d_{i}\neq 0$ and $\sigma_{n}$ is a permutation of the indices of $(x_{1},\dots,x_{n})$. Then, given $i$, $X_{i}=d_{j}X_{j}$ for some $j$. Then $\operatorname{Var}[X_{i}]=d_{j}^{2}\operatorname{Var}[X_{i}]$, and thus for $\Omega_{A}$ to be non-empty, it must be that $d_{j}\in\{-1,1\}$. If $d_{j}=-1$, then for $X_{i}$ with underlying distribution $P$ to lie in $\Omega_{A}$, it must be that $\operatorname{\mathbb{E}}[X_{i}]=-\operatorname{\mathbb{E}}[X_{i}]$ so that $\operatorname{\mathbb{E}}[X_{i}]=0$ and that $X_{i}=-X_{i}$, which is satisfied if and only if $P$ is symmetric around zero. Thus $\Omega_{A}$ equals the distributions that are symmetric around zero. Finally, if $d_{j}=1$, then $X_{i}=X_{i}$ is satisfied by any distribution, so that $\Omega_{A}=\Omega$. Since $\mathbb{G}$ is a collection of linear invertible matrix, $\Omega_{\mathbb{G}}$ is an intersection of these possible sets of distributions, as required. ∎
@@
There exists a large integer $N$ such that $x_{1},y_{1}],\cdots,[x_{m},y_{m}],q\subset[-N+1,N-1]$. Choose $\alpha>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}\mathbb{1}_{[-N-1,-N]}+(1-\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha})\mathbb{1}_{[N,N+1]}$.
@@ B.2.2 Proof strategy
By definition, a function is either an approximate permutation map, not mass-preserving, or mass-preserving but not box-orbit-preserving. If it is an approximate permutation map, Lemma 1 shows it is invariant to all distributions. If it is not, Lemmas 2-4 show it is not invariant to all distributions in $\Omega_{0}$ when it is locally-dense. The result then follows by Proposition 1. ∎
@@ B.2.3 Proof of lemmas
Proof of Lemma 2.
First, for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, we can find a disjoint partition of $f(A)=\bigsqcup_{i=1}^{n!}B_{i}\bigsqcup C$ such that $B_{i}\subset\sigma_{i}(A)$, $C\subset\bar{A}^{c}$. By Lemma 1, we know that an approximate permutation map is also orbit-preserving. Thus, $m(f(A)\setminus\bar{A})=0$ and $m(C)=0$. Let $A_{i}=f^{-1}(B_{i}).$ We would like to show that $\sigma_{i}(A_{i})=B_{i}$ almost everywhere, i.e., $m(\sigma_{i}(A_{i})\triangle B_{i})=0$, where $\sigma_{i}(A_{i})\triangle B_{i}$ denotes the symmetric difference between two sets $A_{i}$ and $B_{i}$. Then, it follows

### v2

@@ 4.2 Arbitrary groups
We now allow $\mathbb{G}$ to be arbitrary. The classical result of Bahadur and Savage (1956) shows that sufficiently rich classes of distributions cannot be tested in a nonparametric setting. In our context, a related conclusion holds under a local version of this richness condition.
Let $\Omega_{\text{cts}}(\mathcal{B})$ be the set of distributions supported on $\mathcal{B}\subseteq\mathbb{R}$. Informally, a null is locally-dense if it can locally replicate any density on finitely many small intervals (up to scale).
@@
A set $\Omega_{0}\subset\Omega_{\text{cts}}(\mathcal{B})$ is locally-dense if there exists $L:\mathbb{N}\to\mathbb{R}_{+}$ such that for any collection of intervals $I_{1},\dots,I_{m}$ with $|I_{i}|<L(m)$ and any density $p$ supported on $\bigcup_{i=1}^{m}I_{i}$, there exist $H\in\Omega_{0}$ with density $h$ and $\alpha\in(0,1]$ such that
$h(x)=\alpha p(x)\quad\text{for all }x\in\bigcup_{i=1}^{m}I_{i}.$
@@
The proof uses Proposition 1 and shows that for any function $f$ with $\Omega_{0}\subseteq\Omega_{f}$, local denseness implies $\Omega_{f}=\Omega$. Intuitively, local denseness ensures that the null can approximate any density on finitely many regions, allowing the invariance to extend to all distributions.
@@
To illustrate the content of this result, consider the null $\operatorname{\mathbb{E}}_{P}[X_{i}]=\beta$. Let $I_{1},\dots,I_{m}$ be intervals with width bounded by $\delta>0$. For any density $p$ supported on these intervals with mean $\gamma_{p}$ and any $\alpha\in(0,1]$, construct a distribution $H$ with density $\alpha p(x)$ on these intervals. Then $H$ has mean
$\alpha\gamma_{p}+(1-\alpha)\operatorname{\mathbb{E}}_{H}[X_{i}\mid X_{i}\notin\cup_{j}I_{j}].$
By placing the remaining mass on two intervals above and below $\beta$, one can choose $\alpha$ and the remaining mass so that the mean equals $\beta$. Thus the null is locally-dense. The following results all make use of similar constructions.
@@
We show in Supplementary Appendix B.3 that Proposition 7 extends to $\Omega_{\text{cts}}([b_{0},b_{1}])$ for interval $b_{0},b_{1}]$ and $\beta$ in the support. The same extension holds for Proposition 8 for any $q\in(b_{0},b_{1})$. We also show that the null hypothesis of fixed variance is locally-dense and therefore does not satisfy the randomization hypothesis.
@@ 5 Conclusion
Randomization tests yield exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. In practice, randomization tests often rely on invariance conditions that are stronger than the null hypothesis of primary interest. This paper developed a framework for characterizing which null hypotheses admit randomization tests with exact finite-sample validity.
Applying our framework to one-sample problems revealed impossibilities and limited possibilities of uses of finite-sample randomization tests. Many standard null hypotheses, including the null of mean zero, do not admit randomization tests with exact finite-sample validity. We further show that, when restricted to linear group actions, one-sample randomization tests can only test membership in the class of symmetric distributions or in subsets of the Gaussian family. These results imply that existing procedures are not overlooking alternative randomization tests that would achieve finite-sample validity.
While testing beyond symmetry or normality requires non-linear transformations, we show that even such transformations cannot be used to test many commonly studied null hypotheses. More broadly, the framework developed here delineates the limits of finite-sample randomization tests and provides constructive guidance for identifying when nulls admit such tests and how to construct them when they exist. Overall, our findings indicate that current practice is not omitting feasible exact procedures and further motivate a focus on asymptotic validity.
@@ Appendix A Examples of randomization tests
    For $L>0$, let $\mathcal{A}\equiv\{-L,\dots,0,\dots,L\}$ and let $\Omega_{0}$ be the set of distributions that are symmetric around 0. Consider any $x,y$ with $x_{j}\neq 0$ for some $j$ and y=xy=-x. We then have that $c(x_{j};x)>c(x_{j};y)$ and that for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(-x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{0}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that send some values of $x\in\mathcal{A}^{n}$ to x-x and otherwise act as the identity function.
@@
    For any $\mathcal{A}$, let $\Omega_{0}$ be the set of distributions where $p(\alpha_{\ell})=p(\alpha_{\ell^{\prime}})$ for some $\alpha_{\ell}\neq\alpha_{\ell^{\prime}}$. Consider any $x,y$ with $x$ containing $\alpha_{\ell}$ but not $\alpha_{\ell^{\prime}}$ and $y_{j}=\alpha_{\ell^{\prime}}\mathbb{1}[x_{j}=\alpha_{\ell}]+\alpha_{\ell}\mathbb{1}[x_{j}=\alpha_{\ell^{\prime}}]+x_{j}\mathbb{1}[x_{j}\neq\alpha_{\ell},x_{j}\neq\alpha_{\ell^{\prime}}]$. We then have that $c(\alpha_{\ell};x)>c(\alpha_{\ell};y)$ and for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{0}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that acts as the identity except for swapping the values of $\alpha_{\ell}$ and $\alpha_{\ell^{\prime}}$.
@@
    Let $\mathbb{G}$ be the orthogonal group of degree $n$. By Maxwell’s Theorem, the set of invariances given $\mathbb{G}$ is $\Omega_{\mathbb{G}}=\{N(0,\sigma^{2}):\mu\in\mathbb{R},\sigma^{2}>0\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian with mean zero.
@@
    In the below proof, we show that $\mathbb{G}$ is a group and that $\Omega_{\mathbb{G}}=\{N(\mu,\sigma^{2}):\mu\in\mathbb{R},\sigma^{2}>0\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian.
@@
Let $X_{1},X_{2},X_{3}$ be i.i.d. draws from $N(\mu,\sigma^{2})$. Let $X=(X_{1},X_{2},X_{3})$ and $V=(\mu,\mu,\mu)$. Because
@@
Inductively, $A^{n}X\stackrel{{\scriptstyle d}}{{=}}X$ for all $n\in\mathbb{Z}$. Thus, $\{N(\mu,\sigma^{2}):\mu\in\mathbb{R},\sigma^{2}>0\}\subset\Omega_{\mathbb{G}}$. By proposition 5, $\{N(\mu,\sigma^{2}):\mu,\sigma\in\operatorname{\mathbb{R}}\}=\Omega_{\mathbb{G}}$. ∎
@@ B.1 All proofs excluding the proof of Proposition 6
$\Omega_{f}=\bigcap_{x\in\mathcal{A}^{n}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})\ \text{s.t.}\ y=f(x)\right\}.$ (10)
Then any $f$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$ must have some $x^{\star}\in\mathcal{A}^{n}$ and $y^{\star}=f(x^{\star})$ such that
@@
Let $x,y\in\mathcal{A}^{n}$ be arbitrary such that $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$ and let $\beta_{i}(x,y)=c(\alpha_{i};x)-c(\alpha_{i};y)$. We know that there exists some $1\leq i\leq K-1$ such that $\alpha_{i}\leq q<\alpha_{i+1}$. One of the following two cases must occur:
@@
Since $\mathbb{G}$ is a group of linear maps, each element is an invertible $n\times n$ matrix $A=\{a_{ij}\}_{ij}$. Suppose there exists $i,i^{\prime}$ and some $j^{\prime}$ such that $a_{ij^{\prime}},a_{i^{\prime}j^{\prime}}\neq 0$. If $\Omega_{A}$ is not empty, then
@@
where $d_{i}\neq 0$ and $\sigma_{n}$ is a permutation of the indices of $(x_{1},\dots,x_{n})$. Then, given $i$, $X_{i}=d_{j}X_{j}$ for some $j$. Then $\operatorname{Var}[X_{i}]=d_{j}^{2}\operatorname{Var}[X_{i}]$, and thus for $\Omega_{A}$ to be non-empty, it must be that $d_{j}\in\{-1,1\}$. If $d_{j}=-1$, then for $X_{i}$ with underlying distribution $P$ to lie in $\Omega_{A}$, it must be that $\operatorname{\mathbb{E}}[X_{i}]=-\operatorname{\mathbb{E}}[X_{i}]$ so that $\operatorname{\mathbb{E}}[X_{i}]=0$ and that $X_{i}=-X_{i}$, which is satisfied if and only if $P$ is symmetric around zero. Thus $\Omega_{A}$ equals the distributions that are symmetric around zero. Finally, if $d_{j}=1$, then $X_{i}=X_{i}$ is satisfied by any distribution, so that $\Omega_{A}=\Omega$. Since $\mathbb{G}$ is a collection of linear invertible matrices, $\Omega_{\mathbb{G}}$ is an intersection of these possible sets of distributions, as required. ∎
@@
There exists a large integer $N$ such that $x_{1},y_{1}],\cdots,[x_{m},y_{m}],q\subset[-N+1,N-1]$. Choose $\alpha>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}\mathbb{1}_{[-N-1,-N]}+(1-\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha})\mathbb{1}_{[N,N+1]}$.
@@ B.2.2 Proof strategy
By definition, a function is either an approximate permutation map, not mass-preserving, or mass-preserving but not box-orbit-preserving. If it is an approximate permutation map, Lemma B.1 shows it is invariant to all distributions. If it is not, Lemmas B.2-B.4 show it is not invariant to all distributions in $\Omega_{0}$ when it is locally-dense. The result then follows by Proposition 1. ∎
@@ B.2.3 Proof of lemmas
Proof of Lemma B.2.
First, for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, we can find a disjoint partition of $f(A)=\bigsqcup_{i=1}^{n!}B_{i}\bigsqcup C$ such that $B_{i}\subset\sigma_{i}(A)$, $C\subset\bar{A}^{c}$. By Lemma B.1, we know that an approximate permutation map is also orbit-preserving. Thus, $m(f(A)\setminus\bar{A})=0$ and $m(C)=0$. Let $A_{i}=f^{-1}(B_{i}).$ We would like to show that $\sigma_{i}(A_{i})=B_{i}$ almost everywhere, i.e., $m(\sigma_{i}(A_{i})\triangle B_{i})=0$, where $\sigma_{i}(A_{i})\triangle B_{i}$ denotes the symmetric difference between two sets $A_{i}$ and $B_{i}$. Then, it follows